Mr. James is 4 times as old as his daughter. The difference in their ages is 36 years. Find their ages.
Answer by macston(5194)

J=Mr. James age; D=daughter's age
J=4D
J-D=36yrs
4D-D=36yrs
3D=36yrs
D=12yrs ANSWER 1: The daughter is 12 years old.
J=4D=4(12yrs)=48yrs ANSWER 2: Mr. James is 48 years old.
CHECK
J-D=36yrs
48yrs-12yrs=36yrs
36yrs=36yrs